First off, we are asked to find the value of x. Then, we are given a triangle. One angle measures 133 degrees. Another is (x+20)degrees. Another is 2x degrees. How do we solve these types of problems? We have 6 different ones. Thanks.
Answer by Earlsdon(6098)

On problems involving triangles and unknown angles, you can use the well-known fact that the sum of the angles in any plane triangle is always 180 degrees.
So, set up the equation to add the given angles:
133 + (x+20) + (2x) = 180 Combine like-terms.
133 + 3x + 20 = 180 Simplify.
153 + 3x = 180 Subtract 153 from both sides.
3x = 27 Finally, divide both sides by 3.
x = 9 degrees.